OpenStudy (anonymous):

For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 2x + y = 8 -x - 3y = -12

OpenStudy (anonymous):

did you solve for x in the second equation? what did you get?

OpenStudy (anonymous):

no i dont understand it

OpenStudy (anonymous):

"isolate x" means to "solve for x"... this basically means to get x only on one side of the equal sign... so in the second equation....

OpenStudy (anonymous):

i know that. its just the equation is difficult.

OpenStudy (anonymous):

\(\large -x-3y=-12 \) add -3y to both sides: \(\large -x-3y\color {red}{+3y}=-12\color {red}{+3y} \) \(\large -x=-12+3y \) now you can multiply "-1" to both sides: \(\large -x\color {red}{\cdot(-1)}=(-12+3y)\color {red}{\cdot(-1)} \) \(\large x=12-3y\)

OpenStudy (anonymous):

thankk youuuuu

OpenStudy (anonymous):

yw...:)